What is the freezing point of a 0.40 molal solution of sucrose (nonelectrolyte) in ethanol? (Kf of ethanol = 1.99 °C/m, normal freezing point of ethanol is -114.1 °C) *
1
Expert's answer
2020-05-28T10:33:51-0400

using formula or depression in freezing point


∆Tf = -Kf*m

Kf is given 1.99 °C/m

molality is 0.40 molal

so ∆Tf=- 1.99*0.40

∆Tf = -0.796°C

T2-T1=-0.796°C

T1 is given -114.1°C

so T2=-118.96°C